# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

### v2

@@ Abstract
While large multi-modal models (LMMs) have exhibited impressive capabilities across diverse tasks, their effectiveness in handling complex tasks has been limited by the prevailing single-step reasoning paradigm. To this end, this paper proposes VoCoT, a multi-step Visually grounded object-centric Chain-of-Thought reasoning framework tailored for inference with LMMs. VoCoT is characterized by two key features: (1) object-centric reasoning paths that revolve around cross-modal shared object-level information, and (2) visually grounded representation of object concepts in a multi-modal interleaved and aligned manner, which effectively bridges the modality gap within LMMs during long-term generation. Additionally, we construct an instruction dataset to facilitate LMMs in adapting to reasoning with VoCoT. By introducing VoCoT into the prevalent open-source LMM architecture, we introduce VolCano. With only 7B parameters and limited input resolution, VolCano demonstrates excellent performance across various scenarios, surpassing SOTA models, including GPT-4V, in tasks requiring complex reasoning. Our code, data and model will be available at https://github.com/RupertLuo/VoCoT.
11footnotetext:  Co-first authorship.  Corresponding author.
@@ 1 Introduction
In recent years, the success of large language models (LLMs) [36; 37; 45] has been gradually extended to the multi-modal domain. By equipping LLM backbones [45; 46; 8; 16] with visual encoders [40] and efficient cross-modal alignment through generative training on image-text data [31; 42; 4; 63], constructed large multi-modal models (LMMs) possess the capabilities to perceive visual signals and engage in dialogue with users in multi-modal contexts [22; 62; 31; 58; 9; 2].
@@
Despite the potential demonstrated by LMMs to serve as unified and versatile foundations for multi-modal information processing, even models like GPT-4V struggle in compositional tasks requiring complex reasoning [55; 53], such as spatial reasoning [10]. This phenomenon indicates a major limitation of current LMMs: the prevailing single-step question-to-answer (Q2A) inference paradigm that directly generates answers based on questions [9; 29]. As illustrated in Figure 1, correctly answering the question relies on analyzing the actions and relationships of multiple objects and reasoning step-by-step, which is almost impossible to accomplish in a single-step prediction. Moreover, the Q2A paradigm obscures the problem-solving process, limiting the interpretability of the LMM outputs. Conversely, the chain-of-thought (CoT) paradigm, which involves multi-step reasoning, has been widely explored in LLMs [19; 51; 49]. The effectiveness of textual CoTs suggests a promising direction for enhancing LMMs.
@@
Figure 1: An example to compare different inference paradigms in LMMs. (a) A visual question that requires complex reasoning. (b) The conceptual object-centric reasoning path constructed to solve the problem. (c) Outputs of GPT-4V and the proposed VolCano. Hallucination is included in the output of GPT-4V. VoCalno performs multi-step reasoning in VoCoT format. Key objects are highlighted and colors indicate the correspondence between object descriptions and the grounded objects in the image. “box” represents the bounding box coordinates of mentioned objects.
@@
However, in complex contexts where multi-modal information coexists, constructing effective multi-step reasoning paths faces several challenges: (1) Difficulty in integrating reasoning anchors within multi-modal contexts. Textual CoTs mainly extract key information from contexts, such as entities, as anchors and conduct multi-step reasoning around these anchors [57]. In multi-modal contexts, the anchor information is further required to be concepts shared between images and texts and establish connections between modalities. Existing work attempts to introduce anchor information in images, such as segmentation maps [54] and dot matrices [20], asking LMMs to utilize this information during inference. However, these methods require strong capabilities of LMMs to understand these information, thus are almost exclusively applicable to GPT-4V. (2) Limited grounding capabilities of LMMs. During the generation process, LMMs may fail to ground textual descriptions to the corresponding visual information, resulting in erroneous information generated. For example, GPT-4V incorrectly ground the target person to the waiter in Figure 1. The risk of producing hallucination [24; 48] hinders the model from performing effective multi-step reasoning.
@@
To address these challenges, we introduce a framework to empower LMMs for effective and reliable multi-step reasoning. We propose VoCoT: Visually grounded object-centric Chain of Thought, a CoT format that is compatible with LMM inference: (1) As illustrated in Figure 1 (a, b), objects serve as fundamental semantic units in both images and text, effectively bridging multi-modal information. Therefore, VoCoT leverages objects as anchors for reasoning. LMMs are encouraged to conduct multi-step analysis on the properties of key objects, as well as the relationships between them, ultimately reaching a conclusion. (2) To ensure the reliability of information in reasoning paths, VoCoT represents each object in a visually grounded format: a tuple of <textual description, coordinates, corresponding visual representations>. Models are required to explicitly ground objects in the image by generating coordinates for them. Visual representations of objects are supplemented to enhance the cross-modal relevance in reasoning paths. This design mimics the reasoning habit of human, where we continuously reference the visual information of an object in the image when we mention it. (3) We propose a RefBind mechanism to efficiently obtain the visual representations of objects without additional computation. Specifically, RefBind indexes the representation of each object from the complete image representation based on its coordinate. Generally, VoCoT constructs multi-modal interleaved reasoning paths where cross-modal aligned anchors are incorporated.
@@
Nevertheless, there is a significant disparity between VoCoT and the formats of existing visual instruction data. To address this, we further construct a dataset, VoCoT-Instruct-80K, to train LMMs for reasoning in the format of VoCoT. VoCoT-Instruct-80K is built on multiple data sources: (1) Reformulating structured reasoning paths from GQA [15]. (2) Supplementing visual question-answer pairs with thought processes. (3) Constructing complex questions and reasoning paths from images annotated with objects. By curating a wide range of data and leveraging assistance of GPT-4V, the presented dataset maintains both diversity and consistency in the desired format.
@@
Based on the introduced VoCoT framework and dataset, we develop VolCano, a Visually grounded multi-modal Chain-of-thought reasoning model. With only 7B parameters and $336\times 336$ input resolution, VolCano surpasses GPT-4V on various benchmarks that require complex reasoning and compositional capabilities.
2 VoCoT: Visually grounded Object-centric Chain-of-Thoughts
@@
In this section, we explain how to enable LMM to perform multi-step reasoning in the format of VoCoT. In Section 2.1, we elaborate on the formulation of VoCoT. Section 2.2 presents how to utilize existing data resources to construct instruction-tuning datasets that conform to the VoCoT format.
2.1 Formulating VoCoT
@@
VoCoT requires LMMs to perform step-by-step reasoning based on the provided context. Following textual CoTs [19; 51; 49], the reasoning logic in VoCoT is primarily expressed in text but is not limited to specific formats. However, there exists a significant gap between multi-modal and text-only contexts. In order to construct effective and reliable reasoning paths in multi-modal contexts, we characterize VoCoT with two features: (1) Object-centric. Objects are the basic semantic units in images and can serve as anchors to establish connections between multi-modal contextual information. Therefore, VoCoTs are required to include important objects, followed by relevant information extraction and analysis. (2) Visually grounded. Key objects included in VoCoT should be represented by tuples of“<text description, coordinates, visual object representation>”, coordinates ground objects in the images while the visual representation enhances the cross-modal relevance in the reasoning paths.
2.2 VoCoT Dataset Construction
Table 1: Examples to illustrate the construction of VoCoT-formatted data from three data sources. Type 1 data are obtained by rules, while Type 2 and Type 3 data are obtained by leveraging GPT-4V.
Type 1 (GQA Source) Rule Input
[Object Info]: [Uncaptioned image]
Shelf: [0.224, 0.219, 0.386, 0.592], Door: [0.394, 0.176, 0.524, 0.645]
[Question]: What is common to the shelf and the door?
[SemanticStr]: select: shelf -> select: door -> common: [0, 1]
[FullAnswer]: The material, both the shelf and the door are wooden.
[Answer]: material.
Type 1 (GQA Source) Response
[Thought]: Find the shelf [0.224, 0.219, 0.386, 0.592]. Find the door [0.394, 0.176, 0.524, 0.645] . The question ask the common attribute of the two objects. The material, both shelf [0.224, 0.219, 0.386, 0.592] and door [0.394, 0.176, 0.524, 0.645] are wooden. So answer is material.
Type 2 (VQA-Based Source) GPT-4V Input
[Object Info]: TV: [0.78, 0.84, 0.97, 0.98], , Tie: [0.72, 0.12, 0.90, 0.60], [Uncaptioned image]
Toothbrush: [0.15, 0.37, 0.56, 0.49], Person: [0.0, 0.13, 0.81, 1.0]
[Question]: What might be the purpose behind the woman’s action?
[Answer]: The purpose behind her actions may be for entertainment.
Type 2 (VQA-Based Source) Response
[Thought]: The woman in the image [0.0, 0.13, 0.81, 1.0] is engaging in an unusual activity by using a toothbrush [0.15, 0.37, 0.56, 0.49] that is significantly larger than a standard one and wearing a tie [0.38, 0.6, 0.59, 0.99] that also appears to be oversized. … It’s plausible that the woman’s actions are meant to entertain or educate an audience.
Type 3 (Image-Only Source) GPT-4V Input
[Object Info]: Ferris wheel: [0.09, 0.40, 0.14, 0.54] , [Uncaptioned image]
Building: [0.40, 0.36, 0.58, 0.53], Statue: [0.72, 0.12, 0.90, 0.60],
Dock: [0.00, 0.56, 0.33, 0.67], Stair: [0.01, 0.69, 1.00, 1.00],
Type 3 (Image-Only Source) Response
[Question]: Where is this location?
[Refering of Thought]: First, The statue [0.72, 0.12, 0.90, 0.60] is
Merlion Statue, a mythical creature with the head of a lion and
the body of a fish, which is a mascot of Singapore. Secondly, the building [0.40, 0.36, 0.58, 0.53] … Based on the distinctive architecture of the landmarks, this image is taken in Singapore.
[Answer]: Singapore.
@@ 2.2 VoCoT Dataset Construction
The Community has witnessed a surge in multi-modal instruction-following datasets, such as LLaVA-Instruct-150K[31], Valley-Instruct-65K[35] and ALLaVA [4]. However, none of these datasets meet the requirements of the VoCoT format, which includes: (1) responses to human instructions with CoT-formatted multi-step reasoning processes, (2) responses with visually grounded object-centric information, i.e., objects with corresponding coordinates. In this section, we introduce the pipeline to construct a VoCoT-formatted dataset from three types of data sources as summarized below.
@@
GQA [15] is a VQA dataset that includes structured information: each image is paired with a scene graph, and a SQL-like reasoning path over the corresponding scene graph is provided for each VQA pair. An example is exhibited in the first part of Table 1. Inspired by Shikra [6], we convert the SQL-like query statements (“[SementicStr]") and answers (“[FullAnswer]") into fluent and complete textual reasoning thoughts with a rule-based approach. Object bounding boxes from the GQA source are also utilized to supplement the anchor object information that appears in the verbalized reasoning thoughts.
@@
Another intuitive way to construct data in VoCoT format is to supplement VQA data with multi-step reasoning processes in the middle of the question-to-answer process. With the assistance of GPT-4V, reasoning thoughts are generated based on images, questions, answers, and object information within the images. The second part of Table 1 provides an example. Furthermore, we control the output format through in-context learning. Specifically, a crafted sample is included in the input context. As overly simple questions may not require complex reasoning, we sample a subset of data from complex reasoning problems in LLaVA-Instruct [30] as the source.
@@
Although the aforementioned two construction methods are effective, the generated data is limited to existing questions. To enhance the richness of questions and reasoning logic, we leverage the powerful generative capabilities of GPT-4V to expand the constructed dataset. As shown in the third part of Table 1, GPT-4V is provided with images and object information and is prompted to generate complex questions, along with VoCoT-formatted reasoning paths and answers. In-context samples are also incorporated to ensure the correct output format. We choose LVIS [12] as the data source due to the diversity of objects included.
Ultimately, we present VoCoT-Instruct-80K, comprising 72K, 6K, and 2K samples constructed from Type 1, Type 2, and Type 3 sources, respectively. More details about the dataset construction process, including the rule-based conversion approach, prompts for GPT-4V, and quality control methods used are provided in Appendix A.1.
@@ 3.1 Architecture
Figure 2: Illustration of the VolCano framework. The input and output are shown below and above the model, respectively. The blue and green rounded rectangles represent textual and visual tokens, respectively. Special tokens “[c]” and “[/c]” denotes the beginning and end of the coordinates (“[coor.]” in the figure). Coordinates are represented in text. In the output, we visualize coordinates by drawing corresponding boxes in the image for a better illustration. RefBind computes the visual representations of objects with the image features and predicted coordinates.
The overall model architecture of VolCano is shown in Figure 2. VolCano is built on top of a causal decoder-only LLM as the backbone. We incorporate a vision transformer (ViT) as the visual encoder to encode image inputs. A two-layer MLP is adopted as the connection module to map the output of the visual encoder into the input space of the language backbone.
@@
VolCano represents image-text data as an interleaved sequence of visual and textual tokens. Text inputs are tokenized and represented using the token embedding layer. Images and objects can appear at any position in the sequence and are represented by visual tokens. Images are encoded by the visual encoder. The resulting 2D feature maps are flattened into a 1D sequence of visual tokens. The connection module is applied to map the visual tokens to the same dimension as the embedded textual tokens.
Following the configuration of VoCoT in Section 2.1, each object is represented by a visually grounded format: “{textual description} [c] {coordinates} [/c] {visual representation}”, e.g., “dog [c] 0.27, 0.08, 0.92, 0.81 [/c] $V_{\textrm{dog}}$”. “[c]” and “[/c]” are special tokens denoting the beginning and ending of coordinates. We use bounding boxes $[x_{\textrm{min}},y_{\textrm{min}},x_{\textrm{max}},y_{\textrm{max}}]$ as coordinates of objects, $x$ and $y$ are normalized between 0 and 1 with respect to the size of the image. Coordinates are treated as text, undergoing tokenization and embedding.
In addition to text and coordinates, visual tokens of objects, such as $V_{\textrm{dog}}$, are supplemented to help the model reference the corresponding visual information in images. The visual tokens are obtained based on the coordinates and image tokens through the RefBind mechanism. The details of RefBind are explained in the next paragraph. Once the end of coordinates token, “[/c]”, is detected in the input or generated sequence, RefBind is activated to obtain the visual object tokens based on the coordinate between “[c]” and “[/c]”. The obtained object tokens are further appended after the “[/c]” token.
Figure 3: Illustration of the RefBind mechanism.
RefBind Mechanism
A straightforward method for representing objects is to crop the corresponding region in the image and encode each sub-image with the visual encoder. However, this method introduces additional computational costs and loses the contextual information of the complete image. For regions with very few pixels, representing objects with sub-images would introduce redundant information. To tackle with above issues, we propose the RefBind mechanism.
RefBind (short for “Reffering Bind”) is conceptually illustrated in Figure 3. Inspired by the RoI-align method in Faster-RCNN [41], given a bounding box and the encoded 2D grid features of the entire image, we can efficiently index the patches in which the target object appears. The features of these patches are flattened into a sequence that represents the object. RefBind relies solely on indexing operations without additional computation. Additionally, the object representation sequence obtained by RefBind inherently preserves contextual information of the whole image.

### v3

@@ Abstract
While large multi-modal models (LMMs) have exhibited impressive capabilities across diverse tasks, their effectiveness in handling complex tasks has been limited by the prevailing single-step reasoning paradigm. To this end, this paper proposes VoCoT, a multi-step Visually-grounded object-centric Chain-of-Thought reasoning framework tailored for inference with LMMs. VoCoT is characterized by two key features: (1) object-centric reasoning paths that revolve around cross-modal shared object-level information, and (2) visually grounded representation of object concepts in a multi-modal interleaved and aligned manner, which effectively bridges the modality gap within LMMs during long-term generation. To adapt LMMs in reasoning with VoCoT, we further construct an instruction-tuning dataset. By combining VoCoT with the prevalent open-source LMM architectures, we develop a VoCoT-based model, VolCano. With only 7B parameters and limited input image resolution, VolCano demonstrates excellent performance across various scenarios. In benchmarks like CLEVR and EmbSpatial, which highly require complex reasoning capabilities, VolCano outperforms SOTA models, including powerful GPT-4V. Related code, models, and datasets are released in https://github.com/RupertLuo/VoCoT.
VoCoT: Unleashing Visually Grounded Multi-Step Reasoning
in Large Multi-Modal Models
Zejun Li1222Equal contribution., Ruipu Luo2222Equal contribution., Jiwen Zhang1, Minghui Qiu2, Xuanjing Huang3, Zhongyu Wei1,4,5111Corresponding authors. 1School of Data Science, Fudan University 2ByteDance 3School of Computer Science, Fudan University 4Research Institute of Intelligent and Complex Systems, Fudan University 5Shanghai Innovation Institute {zejunli20,zywei}@fudan.edu.cn
@@ 1 Introduction
In recent years, the success of large language models (LLMs) OpenAI (2023a, b) has been gradually extended to the multi-modal domain. By equipping LLM backbones (Touvron et al., 2023a, b; Chiang et al., 2023) with visual encoders (Radford et al., 2021) and efficient cross-modal alignment through generative training on image-text data (Liu et al., 2024b; Schuhmann et al., 2021), the constructed large multi-modal models (LMMs) possess the capabilities to perceive visual signals and engage in dialogue with users in multi-modal contexts Liu et al. (2024b); Dai et al. (2023); Bai et al. (2023).
@@
Despite the potential demonstrated by LMMs to serve as unified and versatile foundations, even models like GPT-4V struggle in composite tasks requiring complex analysis Yang et al. (2023b); Wu and Xie (2023), such as spatial reasoning Du et al. (2024). We attribute this phenomenon to a major limitation of current LMMs: the prevailing single-step question-to-answer (Q2A) inference paradigm that directly generates answers based on questions Dai et al. (2023); Liu et al. (2023b). As illustrated in Figure 1, correctly answering the question relies on analyzing the actions and relationships of multiple objects and thinking step-by-step, which is almost impossible to accomplish in a single-step prediction. Moreover, the single-step Q2A paradigm obscures the problem-solving process, limiting the interpretability of the LMM outputs. Conversely, in the language domain, the chain-of-thought (CoT) paradigm, which involves multi-step reasoning, has been widely explored in LLMs Kojima et al. (2022); Wei et al. (2022), indicating a promising way for enhancing LMMs.
@@
Figure 1: An example to compare different inference paradigms in LMMs. (a) A visual question that requires complex reasoning. (b) The conceptual object-centric reasoning path constructed to solve the problem. (c) Outputs of GPT-4V and the proposed VolCano. Hallucination is included in the output of GPT-4V. VoCalno performs multi-step reasoning in the VoCoT format. In the reasoning path, key objects are highlighted and colors indicate the correspondence between object descriptions and the grounded regions in the image. “[box]” represents the coordinates of mentioned objects. Visual representations of objects are omitted for brevity.
@@
However, for complex contexts where multi-modal information coexists, constructing effective multi-step reasoning paths faces several challenges: (1) Difficulty in integrating reasoning anchors within multi-modal contexts. Textual CoTs mainly extract key information from contexts, such as entities, as anchors and conduct multi-step reasoning around these anchors Yao et al. (2024). In multi-modal contexts, the anchor information is further required to be concepts shared between images and texts and establish connections between modalities. Existing works either supplement the image with additional information (such as segmentation maps Yang et al. (2023a) and dot grids Lei et al. (2024)) as anchors, but such information can only be effectively utilized by GPT-4V, or they roughly consider a single region as the anchor through a search-based approach Shao et al. (2024), failing to model complex multi-object interactions. (2) Limited grounding capabilities of LMMs. During the generation process, LMMs may fail to ground textual descriptions to the corresponding visual information, resulting in erroneous information generated. For example, GPT-4V incorrectly ground the target person to the waiter in Figure 1. The risk of hallucination Li et al. (2023d); Wang et al. (2023b) further hinders effective multi-step reasoning.
@@
To address these challenges, we introduce a framework to empower LMMs for effective and reliable multi-step reasoning. We propose VoCoT, Visually grounded object-centric Chain of Thought. VoCoT is a CoT format that is compatible with LMM inference: (1) As illustrated in Figure 1 (a, b), objects serve as fundamental semantic units in both images and text, effectively bridging multi-modal information. Therefore, VoCoT leverages objects as anchors for reasoning. LMMs are encouraged to conduct multi-step analysis on the properties of key objects, as well as the relationships between them, ultimately reaching a conclusion. (2) To ensure the reliability of reasoning paths, VoCoT represents objects in a visually grounded format: a tuple of <textual description, coordinates, corresponding visual representations>. Models are required to explicitly ground objects in images by generating coordinates for them. Visual representations of objects are supplemented to enhance the cross-modal relevance in reasoning paths. This design mimics the habit of human, where we continuously reference the visual information of an object in the image when we mention it. (3) We propose a RefBind mechanism to efficiently obtain the representations of objects without extra computation. Specifically, RefBind indexes the representation of each object from the image representation based on its coordinates. Generally, VoCoT constructs multi-modal interleaved reasoning paths where cross-modal aligned anchors are incorporated as shown in Figure 1 (c).
@@
Nevertheless, there is a significant disparity between VoCoT and the formats of existing visual instruction data. To this end, we further construct a dataset, VoCoT-Instruct-80K, to train LMMs for reasoning in the format of VoCoT. VoCoT-Instruct-80K is built on multiple data sources: (1) Verbalizing structured reasoning paths from GQA Hudson and Manning (2019). (2) Supplementing visual QA pairs with thought processes. (3) Constructing complex questions and reasoning paths from images annotated with objects. By curating a wide range of data and leveraging assistance of GPT-4V, the presented dataset maintains both diversity and consistency in the desired format.
@@
Based on the introduced VoCoT framework and dataset, we develop VolCano, a Visually-grounded multi-modal Chain-of-thought reasoning model. With only 7B parameters and $336^{2}$ input resolution, VolCano excels in various scenarios and even surpasses GPT-4V on benchmarks like CLEVR Johnson et al. (2017) and EmbSpatial Du et al. (2024) that highly require complex reasoning.
2 Visually-grounded Object-centric CoT
@@
In this section, we explain how to enable LMM to perform multi-step reasoning in the format of visually-grounded object-centric chain-of-thought (VoCoT). In Section 2.1, we elaborate on the formulation of VoCoT. In Section 2.2, we present how to transform existing data resources into instruction-tuning datasets aligned with the VoCoT format.
2.1 VoCoT Formulation
@@
VoCoT requires LMMs to perform step-by-step reasoning based on the provided context. Following textual CoTs Kojima et al. (2022); Wei et al. (2022); Wang et al. (2022), the reasoning logic in VoCoT is primarily expressed in text but is not limited to specific formats. However, there exists a significant gap between multi-modal and text-only contexts. In order to construct effective and reliable reasoning paths in multi-modal contexts, we characterize VoCoT with two features: (1) Object-centric. Objects are the basic semantic units in images and can serve as anchors to establish connections between multi-modal contextual information. Therefore, VoCoTs are required to include important objects, followed by relevant information extraction and analysis. (2) Visually-grounded. Key objects included in VoCoT should be represented by tuples of <text description, coordinates, visual object representation>. During inference, LMMs are required to generate both text and coordinates for objects to explicitly ground them within the images. The visual representation of objects further enhances the cross-modal relevance in the reasoning paths. Section 3.1 introduces how to obtain the visual object representations within current LMM frameworks.
2.2 VoCoT-Instruct-80K Dataset
@@
The community has witnessed a surge in multi-modal instruction-following datasets Liu et al. (2024b); Luo et al. (2023); Chen et al. (2024). However, none of these datasets meet the requirements of the VoCoT format, which includes responses to instructions (1) with CoT-formatted multi-step reasoning processes and (2) with visually grounded object-centric information, i.e., objects with corresponding coordinates. In this section, we introduce the pipeline to construct a VoCoT-formatted dataset from three types of existing data sources.
@@
GQA Hudson and Manning (2019) is a VQA dataset that includes structured information: each image is paired with a scene graph, and a SQL-like reasoning path over the scene graph is provided for each VQA pair. An example is shown by the first part of Table 6 in Appendix A.1. Inspired by Shikra Chen et al. (2023b), we use a rule-based method to verbalize the SQL-like statements “[SementicStr]" and answers “[FullAnswer]" into fluent textual thoughts, supplementing objects descriptions with the corresponding coordinates from the scene graph.
@@
Another intuitive way to construct data in VoCoT format is to supplement VQA data with multi-step reasoning processes in the middle of the Q2A process. With the assistance of GPT-4V, reasoning thoughts are generated based on images, questions, answers, and object information within the images. The second part of Table 6 provides an example. Furthermore, we control the output format through in-context learning. Specifically, a crafted sample is included in the input context. As overly simple questions may not require complex reasoning, we sample a subset of data from complex reasoning problems in LLaVA-Instruct (Liu et al., 2024a) as the source.
@@
Although the aforementioned two construction methods are effective, the generated data is limited to existing questions. To enhance the richness of questions and reasoning logic, we leverage GPT-4V to expand the constructed dataset. As illustrated by Table 6 in Appendix A.1, GPT-4V is provided with images and object information and prompted to generate complex questions, along with VoCoT-formatted reasoning paths and answers. In-context samples are also incorporated to ensure the correct output format. We choose LVIS Gupta et al. (2019) as the data source due to the diversity of objects included.
Ultimately, we construct VoCoT-Instruct-80K, comprising 72K, 6K, and 2K samples from data sources of Type 1, 2, and 3, respectively. More details about the construction process, including the rule-based conversion approach, prompts for GPT-4V, in-context samples, and quality control methods used are provided in Appendix A.1.
Figure 2: Illustration of the VolCano framework. The input and output are shown below and above the model, respectively. The blue and green rounded rectangles represent textual and visual tokens, respectively. Special tokens “[c]” and “[/c]” denotes the beginning and end of the coordinates (“[coor.]” in the figure). Coordinates are represented in text. In the output, we visualize coordinates by drawing corresponding boxes in the image for a better illustration. RefBind obtains the representations of objects with the image features and predicted coordinates.
Figure 3: Illustration of the RefBind mechanism.
@@ 3.1 Architecture
As presented in Figure 2, the overall architecture of VolCano mainly follows LLaVA Liu et al. (2023b). VolCano is built on top of a decoder-only LLM as the backbone. We incorporate a vision transformer (ViT) as the visual encoder to encode image inputs. A two-layer MLP is adopted as the connection module to map the output of the visual encoder into the input space of the language backbone.
@@
VolCano represents image-text data as an interleaved sequence of visual and textual tokens. Text inputs are tokenized and represented using the embedding layer. Images and objects can appear at any position in the sequence and are represented by visual tokens. Images are encoded by the ViT. The obtained 2D feature maps are flattened into 1D sequences and further mapped to visual input tokens through the connection module.
Following the configuration of VoCoT, each object is represented by a visually grounded format: “{textual description} [c] {coordinates} [/c] {visual representation}”, e.g., “dog [c] 0.27, 0.08, 0.92, 0.81 [/c] $V_{\textrm{dog}}$”. “[c]” and “[/c]” are special tokens denoting the beginning and ending of coordinates. We use bounding boxes $[x_{\textrm{min}},y_{\textrm{min}},x_{\textrm{max}},y_{\textrm{max}}]$ as coordinates, $x$ and $y$ are normalized between 0 and 1 w.r.t to the image size. Coordinates are treated as text, undergoing tokenization and embedding.
In addition to text and coordinates, visual tokens of objects, such as $V_{\textrm{dog}}$, are supplemented to help the model reference the corresponding visual information in images. The visual tokens of objects are obtained based on the coordinates and image tokens through the RefBind mechanism. Once the end of coordinates token, “[/c]”, is detected in the input or generated, RefBind is activated to obtain the visual object tokens based on the coordinate between “[c]” and “[/c]”. The obtained object tokens are further appended after the “[/c]” token.
RefBind
A straightforward method for representing objects is to crop the corresponding regions in the image and encode them with the ViT. However, this method introduces additional computational costs and loses the contextual information of the complete image. For regions with very few pixels, representing objects with sub-images would introduce redundant information. To tackle with above issues, we propose the RefBind mechanism.
RefBind (short for “Reffering Bind”) is conceptually illustrated in Figure 3. Inspired by the RoI-pooling method in Fast-RCNN Girshick (2015), given a bounding box and the encoded 2D grid features of the entire image, we can efficiently index the patches in which the target object appears. The features of these patches are flattened into a sequence that represents the object. RefBind relies solely on indexing operations without additional computation. Additionally, the object representation obtained by RefBind inherently preserves contextual information within the whole image.